# Ordered random walks and the Airy line ensemble

### v1

@@ 5.2 Preliminary lemmas
for sufficiently large $t$. Note that $h_{[j]}(x)\leq h(x)$ holds for all $x$ by the condition that the random variables $\zeta_{j}$ in the definition of $h$ take values in $[1,\infty)$. Finally we use that $h(x)\leq CV(x)$ since $x\in W_{t,\varepsilon}$ by Proposition 3.
Consider the second term in (5.2). Since $x\in W_{t,\varepsilon}$ we have $V(x)\leq C\Delta(x)$. Therefore
@@
$\leq C\mathbf{E}[(S_{d}(t)-S_{1}(t))^{d^{2}};(S_{d}(j)-S_{1}(j))>t^{\theta}].$
For every $y\geq t^{\theta}$ we have $\mathbf{P}(S_{d}(t)-S_{1}(t)\geq y)\leq e^{-ry}$ for some $r>0$ for $\lvert x_{j}\rvert\ll t^{\theta}$. Therefore
$\mathbf{E}[(S_{d}(t)-S_{1}(t))^{d^{2}};(S_{d}(t)-S_{1}(t))>t^{\theta}]\leq Ct^{d^{2}}e^{-rt^{\theta}}\rightarrow 0$
for $d^{2}\ll t/\log(t)$. Finally consider the third term in (5.2). For some $r_{1}>0$,
$\mathbf{E}\bigg{[}\prod_{i<j}(S_{j}(t)-S_{i}(t)+\eta_{j+1}-\eta_{i});\eta_{d}>t,\mathcal{T}_{t}\bigg{]}$ $\leq\mathbf{E}[(t+\eta_{d})^{d^{2}};\eta_{d}>t]$
$\leq Ct^{d^{2}}e^{-r_{1}t}\rightarrow 0$
for $d^{2}\ll t/\log(t)$.
@@
Let $a<1/2$ and $\theta>1$. There exists $t_{0}>0,\delta>0$ and $c,C>0$ such that uniformly in $d(t)\leq t^{a}$ and $t>t_{0}$ and $\varepsilon^{\prime}<\varepsilon$, for all $x\in W^{d}$ with $\lvert x_{j}\rvert\ll t^{\theta}$,
$\mathbf{E}_{x}[V(S(t));\tau>t,\nu>t^{1-\varepsilon}]$ $\leq CV(x)e^{t^{(4+\delta)a}-ct^{\epsilon^{\prime}}};$
@@
$\mathbf{E}_{x}[\Delta(B(t));\tau>t,\nu>t^{1-\varepsilon}]$ $\leq C\Delta(x)e^{t^{2a}\log(t)-ct^{\epsilon^{\prime}}}.$
@@
$\mathbf{E}_{x}[V(S(t));\tau>t,\nu>t^{1-\varepsilon}]$ $\leq\mathbf{E}_{x}[V(S(t))^{p}]^{1/p}\mathbf{P}_{x}(\tau>t,\nu>t^{1-\varepsilon})^{1/q}.$
The function $V(x)$ is increasing in differences $(x_{j}-x_{i})$. Choose $\delta$ such that $2a<1/2-\delta$. For every $x\in W$, we set
$x_{j}^{*}=x_{j}+(j-1)t^{1/2-\delta},\quad j=1,\ldots,d.$
Since $x^{*}\in W_{t,\delta}$, by Proposition 3,
$V(x)\leq V(x^{*})\leq C\Delta(x^{*})\leq C(x_{d}-x_{1}+dt^{1/2-\delta})^{d^{2}}.$ (16)
Therefore for $a<1/2$
$(\mathbf{E}_{x}[V(S(t))^{p};(S_{d}(t)-S_{1}(t))\leq t^{\theta}])^{1/p}$
$\leq C(\mathbf{E}[(S_{d}(t)-S_{1}(t)+dt^{1/2-\delta})^{pd^{2}};(S_{d}(j)-S_{1}(j))\leq t^{\theta}])^{1/p}$
@@
Then
$\mathbf{E}_{x}[V(S(t))^{p};(S_{d}(t)-S_{1}(t))>t^{\theta}]$
$\leq C\mathbf{E}[(S_{d}(t)-S_{1}(t)+dt^{1/2-\delta})^{pd^{2}};(S_{d}(j)-S_{1}(j))>t^{\theta}]$
$\leq C\exp(pd^{3}t^{-1/2-\delta})\mathbf{E}[(S_{d}(t)-S_{1}(t))^{pd^{2}};(S_{d}(t)-S_{1}(t))>t^{\theta}]$
where the final factor can be bounded as in Lemma 16 using the bound $\lvert x_{j}\rvert\ll t^{\theta}$. We now use Lemma 13 followed by Lemma 14 to establish that
$\mathbf{E}_{x}[V(S(t));\tau>t,\nu>t^{1-\varepsilon}]$ $\leq Ct^{\theta d^{2}}e^{-ct^{\varepsilon^{\prime}}}$
@@
Let $a<1/2$. There exists $t_{0}>0$ such that uniformly in $d(t)\leq t^{a}$ and $t>t_{0}$ and $x\in W^{d}$ for some $\kappa>1/2-\varepsilon/2$ with $\lvert x_{j}\rvert\leq t^{\theta}$ for $\theta<\kappa\wedge 1$
$\mathbf{E}_{x}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},M_{d}(\nu_{t})>t^{\kappa}]\leq CV(x)e^{t^{(4+\delta)a}-t^{2\kappa-1+\varepsilon}}.$
@@
By the harmonicity of $V$,
$\mathbf{E}_{x}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},M_{d}(\nu_{t})>t^{\kappa}]$
$\leq\mathbf{E}_{x}[V(S(\nu_{t}\wedge t^{1-\varepsilon}));\tau>\nu_{t}\wedge t^{1-\varepsilon},M_{d}(\nu_{t}\wedge t^{1-\varepsilon})>t^{\kappa}]$
$=\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},M_{d}(\nu_{t}\wedge t^{1-\varepsilon})>t^{\kappa}]$
$\leq\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},M_{d}(t^{1-\varepsilon})>t^{\kappa}].$
For $\delta$ satisfying $2a<1/2-\delta$ we use (16) to obtain that
$\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t]$
$\leq C\mathbf{E}_{x}\bigg{[}\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j)+dt^{1/2-\delta})^{d^{2}};\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t\bigg{]}$
$\leq C\exp(d^{3}t^{-1/2-\delta})\mathbf{E}_{x}\bigg{[}\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))^{d^{2}};\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t\bigg{]}.$
@@
and for every $u\geq t$ we have $\mathbf{P}_{x_{d}-x_{1}}(S_{d}(t^{1-\varepsilon})-S_{1}(t^{1-\varepsilon})\geq u)\leq e^{-ru}$ for some $r>0$ and all $\lvert x_{d}-x_{1}\rvert\leq t^{\theta}$. Therefore
$\mathbf{E}_{x}\bigg{[}\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))^{d^{2}};\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t\bigg{]}\leq Ct^{d^{2}}e^{-rt}.$
@@
$\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t]$
$\leq C\exp(d^{3}t^{-1/2-\delta})t^{d^{2}}e^{-rt}\rightarrow 0$
for $d^{2}\ll t/\log t$. We next analyse the expectation over the event $\mathcal{E}_{t}=\{\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))\leq t\}$. On this set we shall use the bound $V(x)\leq h(x)$. We first estimate that for some $r_{1}>0$,
$\mathbf{E}\bigg{[}\prod_{i<j}(S_{j}(t^{1-\varepsilon})-S_{i}(t^{1-\varepsilon})+\eta_{j+1}-\eta_{i});\eta_{d}>t,\mathcal{E}_{t}\bigg{]}$ $\leq\mathbf{E}[(t+\eta_{d})^{d^{2}};\eta_{d}>t]$
$\leq C(2t)^{d^{2}}e^{-r_{1}t}\rightarrow 0$
for $d^{2}\ll t/\log(t)$. If $\eta_{d}\leq t$ and $S_{d}(t^{1-\varepsilon})-S_{1}(t^{1-\varepsilon})\leq t$ then
$h(S(t^{1-\varepsilon}))\leq(2t)^{d}h_{[d]}(S_{[d]}(t^{1-\varepsilon})).$
As a result we have
$\mathbf{E}_{x}[h(S(t^{1-\varepsilon});\tau>t^{1-\varepsilon},M_{d}>t^{\kappa},\mathcal{E}_{t},\eta_{d}\leq t]$
$\leq(2t)^{d}\mathbf{P}_{x_{d}}(M_{d}(t^{1-\varepsilon})>t^{\kappa})\mathbf{E}_{x_{[d]}}[h_{[d]}(S_{[d]}(t^{1-\varepsilon}));\tau_{[d]}>t^{1-\varepsilon}]$
$\leq h_{[d]}(x_{[d]})(2t)^{d}\mathbf{P}_{x}(M_{d}(t^{1-\varepsilon})>t^{\kappa}).$
By an exponential bound for $M_{d}$, for some $r_{2}>0$ and all $\lvert x_{d}\rvert\leq t^{\theta},$
$\mathbf{P}_{x_{d}}(M_{d}(t^{1-\varepsilon})>t^{\kappa})\leq\exp(-r_{2}t^{2\kappa-1+\varepsilon}).$
As a result we have
$\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},M_{d}(t^{1-\varepsilon})>t^{\kappa},\mathcal{E}_{t},\eta_{d}\leq t]\\ \leq Ch_{[d]}(x_{[d]})e^{t^{a}\log(2t)-t^{2\kappa-1+\varepsilon}}.$
Then
$\mathbf{E}_{x}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},M_{d}(\nu_{t})>t^{\kappa}]\leq Ch(x)e^{t^{a}\log(2t)-t^{2\kappa-1-\varepsilon}}.$
Finally we use that $h(x)\leq t^{\theta d^{2}}$ using $\lvert x_{j}\rvert\leq t^{\theta}$ and $V(x)\geq C_{1}\exp(-cd^{4+\delta})$ by Lemma 14. ∎
@@
Uniformly over all $y\in W^{d}$ satisfying $-\theta\leq y_{1}\leq\ldots\leq y_{d}\leq\theta$ we have for uniformly continuous $f$ depending on finitely many co-ordinates
$\mathbf{E}_{y}^{(\Delta)}[f(B(s):0\leq s)]-\mathbf{E}_{0}^{(\Delta)}[f(B(s):0\leq s)]\leq C\theta.$
We prove this statement by using the following monotonicity property of non-intersecting Brownian motions, that also holds for Dyson Brownian motion with general $\beta$.
@@
Suppose $(Y_{1},\ldots,Y_{d})$ and $(Z_{1},\ldots,Z_{d})$ satisfy Dyson Brownian motion run with the same Brownian motions. Suppose that $Y_{j}(0)-Y_{i}(0)\geq Z_{j}(0)-Z_{i}(0)$ for all $1\leq i<j\leq d$. Then for all $t\geq 0$ we have that $Y_{j}(t)-Y_{i}(t)\geq Z_{j}(t)-Z_{i}(t)$ for all $1\leq i<j\leq d.$
@@
We have that $-\theta\leq y_{1}\leq\ldots\leq y_{d}\leq\theta$ so we can run three systems of Dyson Brownian motions started with all particles at $-\theta$, particles at $y$ and with all particles started from $\theta.$ Lemma 20 shows that the spacings remain larger for all time when particles are started at $y$ than for the other two systems. The initial extremal particles are also ordered and hence the particle positions are ordered in the following way. Let $Z$ and $Y$ be Dyson Brownian motions started at 00 and $y$ respectively and run with the same Brownian motions. We have for $j=1,\ldots,d$ and all $s\geq 0$,
@@
$\mathbf{E}_{y}[f(B)]-\mathbf{E}_{0}[f(B)]\leq C\theta.\qed$

### v2

@@ 5.2 Preliminary lemmas
for sufficiently large $t$. Note that $h_{[j]}(x)\leq h(x)$ holds for all $x$ by the condition that the random variables $\zeta_{j}$ in the definition of $h$ take values in $1,\infty)$. Finally we use that $h(x)\leq CV(x)$ since $x\in W_{t,\varepsilon}$ by Proposition 3 and Lemma 10.
Consider the second term in (5.2). Since $x\in W_{t,\varepsilon}$ we have $V(x)\leq h(x)\leq\ C\Delta(x)$ by Proposition 8 and Lemma 10 after which we can use the bound $\Delta(x)\leq(x_{d}-x_{1})^{d^{2}}$ for $x\in W^{d}$. Therefore,
@@
$\leq C\mathbf{E}_{x}[(S_{d}(t)-S_{1}(t))^{d^{2}};(S_{d}(t)-S_{1}(t))>t^{\theta}].$
For every $y\geq t^{\theta}$ we have $\mathbf{P}(S_{d}(t)-S_{1}(t)\geq y)\leq e^{-ry}$ for some $r>0$ for $\lvert x_{j}\rvert\leq t^{\theta}/\log(t)$. Therefore,
$\mathbf{E}_{x}[(S_{d}(t)-S_{1}(t))^{d^{2}};(S_{d}(t)-S_{1}(t))>t^{\theta}]\leq Ct^{\theta d^{2}}e^{-rt^{\theta}}\rightarrow 0$
as $a<1/2<\theta/2$. Finally consider the third term in (5.2). For some $r_{1}>0$,
$\mathbf{E}_{x}\bigg[\prod_{i<j}(S_{j}(t)-S_{i}(t)+\eta_{j}-\eta_{i});\eta_{d}>t^{\theta},\mathcal{T}_{t}\bigg]$ $\leq\mathbf{E}[(t+\eta_{d})^{d^{2}};\eta_{d}>t^{\theta}]$
$\leq Ct^{\theta d^{2}}e^{-r_{1}t^{\theta}}\rightarrow 0$
since $a<1/2<\theta/2$. The condition $\theta>\beta/2$ ensures the second and third terms in (5.2) do not contribute at the leading order, which allows a simpler form for the right hand side in the final statement in (i).
@@
Let $a<1/4$, $\theta>1$ and $\varepsilon\in(0,1/2)$. There exists $t_{0}>0,\delta>0$ and $c,C>0$ such that uniformly in $d(t)\leq t^{a}$ and $t>t_{0}$ and $\varepsilon^{\prime}<\varepsilon$, for all $x\in W^{d}$ with $\lvert x_{j}\rvert\leq t^{\theta}/\log(t)$,
$\mathbf{E}_{x}[V(S(t));\tau>t,\nu_{t}>t^{1-\varepsilon}]$ $\leq CV(x)e^{t^{(4+\delta)a}-ct^{\epsilon^{\prime}}};$
@@
$\mathbf{E}_{x}[\Delta(B(t));\tau>t,\nu_{t}>t^{1-\varepsilon}]$ $\leq C\Delta(x)e^{\theta t^{2a}\log(t)-ct^{\epsilon^{\prime}}}.$
@@
$\mathbf{E}_{x}[V(S(t));\tau>t,\nu_{t}>t^{1-\varepsilon}]$
$\leq\mathbf{E}_{x}[V(S(t))^{p};\tau>t]^{1/p}\mathbf{P}_{x}(\tau>t,\nu_{t}>t^{1-\varepsilon})^{1/q}.$ (17)
The function $V(x)$ is increasing in differences $(x_{j}-x_{i})$ by Proposition 4(b) of [21]. The intuition is that $\mathbf{P}_{x}(\tau>n)\sim\mathcal{X}V(x)n^{-d(d-1)/4}$ as $n\rightarrow\infty$ from [21]. As $\tau$ is increasing in differences $(x_{j}-x_{i})$, then it is possible to conclude that $V(x)$ is as well. Although [21] considers discrete-time random walks, the argument can also be applied to continuous-time random walks.
Choose $\delta^{\prime}>0$ such that $2a<1/2-\delta^{\prime}$. For every $x\in W^{d}$, we set
$x_{j}^{*}=x_{j}+(j-1)t^{1/2-\delta^{\prime}},\quad j=1,\ldots,d.$
Since $x^{*}\in W_{t,\delta^{\prime}}$, by Proposition 3,
$V(x)\leq V(x^{*})\leq C\Delta(x^{*})\leq C(x_{d}-x_{1}+dt^{1/2-\delta^{\prime}})^{d^{2}}.$ (18)
Therefore, for $a<1/2$,
$(\mathbf{E}_{x}[V(S(t))^{p};(S_{d}(t)-S_{1}(t))\leq t^{\theta},\tau>t])^{1/p}$
$\leq C(\mathbf{E}[(S_{d}(t)-S_{1}(t)+dt^{1/2-\delta^{\prime}})^{pd^{2}};0\leq(S_{d}(j)-S_{1}(j))\leq t^{\theta}])^{1/p}$
@@
Then,
$\mathbf{E}_{x}[V(S(t))^{p};(S_{d}(t)-S_{1}(t))>t^{\theta},\tau>t]$
$\leq C\mathbf{E}[(S_{d}(t)-S_{1}(t)+dt^{1/2-\delta^{\prime}})^{pd^{2}};(S_{d}(t)-S_{1}(t))>t^{\theta}]$
$\leq C\exp(pd^{3}t^{1/2-\delta^{\prime}-\theta})\mathbf{E}[(S_{d}(t)-S_{1}(t))^{pd^{2}};(S_{d}(t)-S_{1}(t))>t^{\theta}]$
where the final factor can be bounded as in Lemma 16 using the bound $\lvert x_{j}\rvert\leq t^{\theta}/\log(t)$. We now apply Lemma 13, taking $\varepsilon^{\prime}=\delta$, to the second factor in (5.2) followed by Lemma 14 to establish that
$\mathbf{E}_{x}[V(S(t));\tau>t,\nu_{t}>t^{1-\varepsilon}]$ $\leq Ct^{\theta d^{2}}e^{-ct^{\varepsilon^{\prime}}}$
@@
Let $a<1/4$, $\delta>0$, $\varepsilon\in(0,1)$, $\kappa>1/2-\varepsilon/2$ and $\theta>\max(1,2\kappa-1+\varepsilon)$. There exists $t_{0}>0$ such that uniformly in $d(t)\leq t^{a}$, $t>t_{0}$ and $x\in W^{d}$, with $\lvert x_{d}-x_{1}\rvert\leq t^{\theta}/\log(t)$,
$\mathbf{E}_{x}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},\widehat{M}(\nu_{t})>t^{\kappa}]\leq CV(x)e^{t^{(4+\delta)a}-t^{2\kappa-1+\varepsilon}}.$
@@
Firstly,
$\mathbf{E}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},\widehat{M}(\nu_{t})>t^{\kappa}]$
$\leq\sum_{i=1}^{d}\mathbf{E}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},M_{i}(\nu_{t})>t^{\kappa}].$
For each $i=1,\ldots,d$, by the harmonicity of $V$,
$\mathbf{E}_{x}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},M_{i}(\nu_{t})>t^{\kappa}]$
$\leq\mathbf{E}_{x}[V(S(\nu_{t}\wedge t^{1-\varepsilon}));\tau>\nu_{t}\wedge t^{1-\varepsilon},M_{i}(\nu_{t}\wedge t^{1-\varepsilon})>t^{\kappa}]$
$=\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},M_{i}(\nu_{t}\wedge t^{1-\varepsilon})>t^{\kappa}]$
$\leq\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},M_{i}(t^{1-\varepsilon})>t^{\kappa}].$
For $\delta^{\prime}>0$ satisfying $2a<1/2-\delta^{\prime}$ we use (18) to obtain that
$\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t^{\theta}]$
$\leq C\mathbf{E}_{x}\bigg[\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j)+dt^{1/2-\delta^{\prime}})^{d^{2}};\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t^{\theta}\bigg]$
$\leq C\exp(d^{3}t^{1/2-\delta^{\prime}-\theta})\mathbf{E}_{x}\bigg[\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))^{d^{2}};\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t^{\theta}\bigg].$
@@
and for every $u\geq t$ we have $\mathbf{P}_{x_{d}-x_{1}}(S_{d}(t^{1-\varepsilon})-S_{1}(t^{1-\varepsilon})\geq u)\leq e^{-ru}$ for some $r>0$ and all $\lvert x_{d}-x_{1}\rvert\leq t^{\theta}/\log(t)$. Therefore,
$\mathbf{E}_{x}\bigg[\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))^{d^{2}};\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t^{\theta}\bigg]\leq Ct^{\theta d^{2}}e^{-rt^{\theta}}.$
@@
$\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))>t^{\theta}]$
$\leq C\exp(d^{3}t^{1/2-\delta^{\prime}-\theta})t^{\theta d^{2}}e^{-rt^{\theta}}\rightarrow 0$ (19)
for $a<1/2$. We next analyse the expectation over the event $\mathcal{E}_{t}=\{\max_{j\leq t^{1-\varepsilon}}(S_{d}(j)-S_{1}(j))\leq t^{\theta}\}$. On this set we shall use the bound $V(x)\leq h(x)$. We first estimate that for some $r_{1}>0$,
$\mathbf{E}\bigg[\prod_{i<j}(S_{j}(t^{1-\varepsilon})-S_{i}(t^{1-\varepsilon})+\eta_{j}-\eta_{i});\eta_{d}>t^{\theta},\mathcal{E}_{t}\bigg]$
$\qquad\leq\mathbf{E}[(t^{\theta}+\eta_{d})^{d^{2}};\eta_{d}>t^{\theta}]\leq C(2t)^{\theta d^{2}}e^{-r_{1}t^{\theta}}\rightarrow 0$ (20)
since $a<1/2$. If $\eta_{d}\leq t^{\theta}$ and $S_{d}(t^{1-\varepsilon})-S_{1}(t^{1-\varepsilon})\leq t^{\theta}$ then
$h(S(t^{1-\varepsilon}))\leq(2t)^{\theta d}h_{[i]}(S_{[i]}(t^{1-\varepsilon})).$
As a result, we have
$\mathbf{E}_{x}[h(S(t^{1-\varepsilon});\tau>t^{1-\varepsilon},M_{i}(t^{1-\varepsilon})>t^{\kappa},\mathcal{E}_{t},\eta_{d}\leq t^{\theta}]$
$\leq(2t)^{\theta d}\mathbf{P}_{x_{i}}(M_{i}(t^{1-\varepsilon})>t^{\kappa})\mathbf{E}_{x_{[i]}}[h_{[i]}(S_{[i]}(t^{1-\varepsilon}));\tau_{[i]}>t^{1-\varepsilon}]$
$\leq h_{[i]}(x_{[i]})(2t)^{\theta d}\mathbf{P}_{x_{i}}(M_{i}(t^{1-\varepsilon})>t^{\kappa}).$
By an exponential bound for $M_{i}$, for some $r_{2}>0$,
$\mathbf{P}_{x_{i}}(M_{i}(t^{1-\varepsilon})>t^{\kappa})\leq\exp(-r_{2}t^{2\kappa-1+\varepsilon}).$
The same exponent holds for all $i=1,\ldots,d.$ As a result we have
$\mathbf{E}_{x}[V(S(t^{1-\varepsilon}));\tau>t^{1-\varepsilon},M_{i}(t^{1-\varepsilon})>t^{\kappa},\mathcal{E}_{t},\eta_{d}\leq t]\\ \leq Ch_{[i]}(x_{[i]})e^{\theta t^{a}\log(2t)-t^{2\kappa-1+\varepsilon}}.$
Then, in combination with (5.2),
$\mathbf{E}_{x}[V(S(\nu_{t}));\tau>\nu_{t},\nu_{t}\leq t^{1-\varepsilon},M_{i}(\nu_{t})>t^{\kappa},\mathcal{E}_{t}]\\ \leq Ch(x)e^{\theta t^{a}\log(2t)-t^{2\kappa-1+\varepsilon}}+C(2t)^{\theta d^{2}}e^{-r_{1}t^{\theta}}.$
We use that $h(x)\leq C(2t)^{\theta d^{2}}$ when $\lvert x_{d}-x_{1}\rvert\leq t^{\theta}/\log(t)$ which follows from
$h(x)$ $=\mathbf{E}\big[\prod_{i<j}(x_{j}-x_{i}+\eta_{j}-\eta_{i});\eta_{d}\leq t^{\theta}\big]+\mathbf{E}\big[\prod_{i<j}(x_{j}-x_{i}+\eta_{j}-\eta_{i});\eta_{d}>t^{\theta}\big]$
$\leq(2t)^{\theta d^{2}}+\mathbf{E}[(t^{\theta}+\eta_{d})^{d^{2}};\eta_{d}>t^{\theta}]$
where the second term can be bounded as in (5.2). We combine this with the bound (5.2) on the event $\mathcal{E}_{t}^{c}.$ Finally, we use that $V(x)\geq C_{1}\exp(-cd^{4+\delta})$ by Lemma 14. ∎
@@
Let $\theta>0$. Uniformly over all $y,z\in W^{d}$ satisfying $\lvert z_{j}-y_{j}\rvert\leq\theta$ for $j=1,\ldots,d$ we have for uniformly continuous $f:C(k,[-L,L])\rightarrow[0,1]$,
$\mathbf{E}_{y}^{(\Delta)}[f(B(s):0\leq s)]-\mathbf{E}_{z}^{(\Delta)}[f(B(s):0\leq s)]\leq C\theta.$
We prove this statement by using the following monotonicity property of non-intersecting Brownian motions. This comes from the representation of non-intersecting Brownian motions as stochastic differential equations, see [29]. We use Dyson Brownian motion to refer to non-intersecting Brownian motions when we use the SDE representation. Note that Dyson Brownian motion has a strong solution, for example see [3]. This representation allows us to couple Dyson Brownian motion with different initial conditions run with the same Brownian motion. In particular, let $(Y_{1}(0),\ldots,Y_{d}(0))=(y_{1},\ldots,y_{d})\in W^{d}$ and $(Z_{1}(0),\ldots,Z_{d}(0))=(z_{1},\ldots,z_{d})\in W^{d}$. For $j=1,\ldots,d$ let
$dY_{j}(t)=\sum_{1\leq i\leq d,i\neq j}\frac{dt}{Y_{j}(t)-Y_{i}(t)}+dB_{j}(t)$
$dZ_{j}(t)=\sum_{1\leq i\leq d,i\neq j}\frac{dt}{Z_{j}(t)-Z_{i}(t)}+dB_{j}(t)$
where we use the same Brownian motions $B_{1},\ldots,B_{d}$ in both systems. The strong solution can be extended to the case when the initial conditions are in the closure of $W^{d}$, see Proposition 4.3.5 of [3].
We use Lemma 2.8 of [1] which we state here. Part (i) also appears as Lemma 4.3.6 of [3] (with strict rather than weak inequalities for the initial condition).
@@
Suppose $(Y_{1},\ldots,Y_{d})$ and $(Z_{1},\ldots,Z_{d})$ are coupled Dyson Brownian motions run with the same Brownian motions as described above started from $y_{1}\leq\ldots\leq y_{d}$ and $z_{1}\leq\ldots\leq z_{d}$ respectively.
  1. (i)
    Suppose that $y_{j}\geq z_{j}$ for all $j=1,\ldots,d$. Then, for all $t\geq 0$ we have that $Y_{j}(t)\geq Z_{j}(t)$ for all $j=1,\ldots,d$.
  2. (ii)
    Suppose that $y_{j}-y_{i}\geq z_{j}-z_{i}$ for all $1\leq i<j\leq d$. Then, for all $t\geq 0$ we have that $Y_{j}(t)-Y_{i}(t)\geq Z_{j}(t)-Z_{i}(t)$ for all $1\leq i<j\leq d.$
@@
We have that $z_{j}-\theta\leq y_{j}\leq z_{j}+\theta$ for $j=1,\ldots,d.$ Let $Z$ and $Y$ be coupled systems of Dyson Brownian motion started at $z$ and $y$ respectively and run with the same Brownian motions. We can also consider the vertical translations given by $Z$ run with the same Brownian motions started from $(z_{1}-\theta,\ldots,z_{d}-\theta)$ and $(z_{1}+\theta,\ldots,z_{d}+\theta)$ respectively. We have for $j=1,\ldots,d$ and all $s\geq 0$ by Lemma 20 part (i),
@@
$\mathbf{E}_{y}[f(B)]-\mathbf{E}_{z}[f(B)]\leq C\theta.\qed$
Alternatively, in the case $z=0$, Lemma 19 could have been established by combining part (ii) of Lemma 20 with the fact that the initial extremal particles are ordered.